Calculate the probability of the following events: 

C = the sum of the digits is less than or equal to 6 

D = the sum of the digits is greater than or equal to 7

 P(C)

 P(D)

 P(C or D)

 P(C and D) 


2 Dice Sample Space 







1,11,21,31,41,51,6
2,12,22,32,42,52,6
3,13,23,33,43,53,6
4,14,24,34,44,54,6
5,15,25,35,45,55,6
6,16,26,36,46,56,6
Answer #1

(a)

P(C) = P(Sum of the digits \leq 6) = 15/36 = 0.4167

So,

Answer is:

0.4167

(b)

P(D) = P(Sum of the digits \geq 7) = 21/36 = 0.5833

So,

Answer is:

0.5833

(c)

P(C or D) = P(Sum of digits \leq 6 OR Sum of digits \geq 7) = 1

So,

Answer is:

1

(d)

P(C and D) = P(Sum of digits \leq 6 AND Sum of digits \geq 7)= 0

So,

Answer is:

0